# A diverse, overlooked population of Type Ia supernovae exhibiting mid-infrared signatures of delayed circumstellar interaction

### v1

@@ Abstract
Type Ia supernovae arise from the thermonuclear explosions of white dwarfs in multiple star systems. A rare sub-class of SNe Ia exhibit signatures of interaction with circumstellar material (CSM), allowing for direct constraints on companion material. While most known events show evidence for dense nearby CSM identified via peak-light spectroscopy (as SNe Ia-CSM), targeted late-time searches have revealed a handful of cases exhibiting delayed CSM interaction with detached shells. Here, we present the first all-sky search for late CSM interaction in SNe Ia using a new image-subtraction pipeline for mid-infrared data from the NEOWISE space telescope. Analyzing a sample of SNe Ia, we report evidence for late-time mid-infrared brightening in six previously overlooked events spanning sub-types SNe Iax, normal SNe Ia, SNe Ia-91T and super-Chandra SNe Ia. Our systematic search doubles the known sample, and suggests that % of SNe Ia exhibit mid-infrared signatures of delayed CSM interaction. The mid-infrared light curves ubiquitously indicate the presence of multiple (or extended) detached CSM shells located at cm, containing of dust, with some sources showing evidence for new dust formation, likely within the cold, dense shell of the ejecta. We do not detect interaction signatures in spectroscopic and radio follow-up; however, the limits are largely consistent with previously confirmed events given the sensitivity and observation phase. Our results highlight that CSM interaction is more prevalent than previously estimated from optical and ultraviolet searches, and that mid-infrared synoptic surveys provide a unique window into this phenomenon.
@@ 1 Introduction
Supporting the diverse nature of progenitors, a rare sub-class of SNe Ia—known as SNe Ia-CSM ( of SNe Ia; Silverman et al. 2013a; Sharma et al. 2023)—exhibit strong interaction of the expanding ejecta111We note that interaction with very nearby CSM can also be detected as early-time excesses in optical light curves or high velocity features (e.g., Noebauer et al. 2016; Mulligan & Wheeler 2017; Moriya et al. 2023); however, we do not discuss them here since existing observations do not conclusively rule out other possibilities (e.g. Childress et al. 2014; Deckers et al. 2022). with dense nearby circumstellar material (CSM). These events are typically identified via strong narrow emission lines (commonly H) in spectroscopy acquired at or near peak light (Sharma et al., 2023), similar to Type IIn supernovae (Smith, 2017). This H-rich material has been ubiquitously attributed to mass supplied by a non-degenerate companion, allowing for direct constraints on its composition and configuration. The proximity ( cm) and large mass ( M⊙) of the CSM have been explained via binary progenitors such as a WD merging with the core of an asymptotic giant branch (AGB) star inside (or shortly after the ejection) of the common envelope (i.e. the core-degenerate scenario; Livio & Riess 2003; Kashi & Soker 2011) or even as a result of dynamical mass transfer from a main sequence star (Han & Podsiadlowski, 2006). Since systematic searches (Silverman et al., 2013a; Sharma et al., 2023) have relied on peak-light spectroscopy or peculiar undulations in early-time light curves to identify candidates for SNe Ia-CSM, these are biased against SNe Ia with distant CSM shells where signatures of interaction would only be detectable at later phases (Sharma et al., 2023; Phillips et al., 2024).
@@ 2.1 MIR behavior of known SNe Ia with late CSM interaction
Using data from the infrared Spitzer Space Telescope (Werner et al., 2004), Fox & Filippenko (2013) reported luminous, brightening late-time MIR emission in SNe Ia-CSM 2002ic and 2005gj, suggesting renewed shock interaction with CSM shells. Subsequently, Szalai et al. (2019), Szalai et al. (2021) and Sharma et al. (2023) also confirmed consistently bright late-time MIR emission in larger samples of SNe Ia-CSM based on serendipitous Spitzer and NEOWISE observations, largely consistent with pre-existing dust heated by ongoing CSM interaction. A similar, luminous MIR excess was also reported in the case of the nearby SN Iax 2014dt (Fox et al., 2016), although its origin has been debated (as arising from circumstellar dust or possibly a bound remnant; Foley et al. 2016). Using a comprehensive spectral sequence and NEOWISE observations, Wang et al. (2024) provided evidence for multiple dust shells as well as new late-time dust formation in the SN Ia-CSM 2018evt. To understand if MIR searches can be fruitful to look for delayed CSM interaction in SNe Ia, we first investigate previously known events in NEOWISE data. The NEOWISE survey’s all-sky coverage, MIR photometric bands (m, m), long baseline ( yr), regular cadence ( month), and sensitivity ( AB mag) are ideal attributes for such a search. When combined with a new difference photometry pipeline (De et al., in prep) with unWISE coadded images (Lang, 2014; Meisner et al., 2018) and a ZOGY-based subtraction algorithm (Zackay et al., 2016; De et al., 2020) adept at recovering faint transients on bright hosts, this MIR dataset has already proven to be powerful for core-collapse SNe (Myers et al., 2024).
NEOWISE observations of PTF11kx begin yrs after peak light, exhibiting luminous emission detected for two more years (also previously reported with Spitzer data in Graham et al. 2017). Similarly, the NEOWISE analysis of SN 2015cp reveals a long-lasting MIR light curve333Thévenot et al. (2021a) previously also reported the NEOWISE detection of SN 2015cp.. We show a comparison of these light curves with SNe Ia-CSM in Figure 1, and their individual light curves in Figure 5. Kool et al. (2023) reported luminous MIR emission from SN 2020eyj that exhibited delayed interaction with a He-rich CSM shell. In contrast, NEOWISE light curves of the sub-luminous SNe Ia which show late-time H emission (SNe 2016jae, 2018cqj, and 2018fhw), as well as the candidates identified from late-time optical photometry (Terwel et al., 2024) (SNe 2018grt, 2019ldf, and 2020tfc) do not show luminous MIR emission comparable to the other events. We discuss possible reasons in Section 3.3. As shown in Figure 1, SNe Ia-CSM and SNe Ia with delayed CSM interaction uniformly show luminous late-time MIR emission, distinct from normal SNe Ia, making MIR searches a promising avenue for identifying this phenomenon.
@@ 2.2 Systematic search in NEOWISE for missed events
We produced NEOWISE MIR difference imaging light curves for Type Ia SNe from 2011 to 2022444The 2022 cutoff is governed by the availability of unWISE data products. reported to the IAU Transient Name Server555https://www.wis-tns.org. These light curves were visually inspected to search for late-time MIR emission associated with the SN. We filter the light curves by i) rejecting candidates with detections prior to the optical onset of the SN (to remove contamination from active galactic nuclei activity) or obvious subtraction artifacts; and ii) requiring at least two epochs of NEOWISE detections ( yr) and rising emission in at least one band, to conclusively rule out excess MIR emission arising purely from new dust formation or a SN light echo, and to eliminate detections deemed to only be associated with photospheric emission from the optical SN. Most of the identified candidates were known SNe Ia-CSM; we do not discuss them here as their light curves have been presented in Sharma et al. (2023). Table 1 summarizes the cuts at each filtering stage. Notably, these selection criteria resulted in the identification of six candidates which exhibit luminous late-time MIR emission despite not being previously associated with CSM interaction.
@@
The candidates span multiple Type Ia subclasses: SN 2017fra (originally classified as a SN Ia-91T), SN 2017hyn (Ia-91T), SN 2020mvp (Ia), SN 2020yex (Ia-91T), SN 2021jun (Iax), and SN 2022bbt (super-Chandra Ia)666Thévenot et al. (2021a) and Thévenot et al. (2021b) have previously noted MIR detections for SN 2017hyn, SN 2017fra, and SN 2020yex; Jiang et al. (2021) previously analyzed NEOWISE observations of SN 2017fra in their search for MIR flares in nearby galaxies.. We further describe each of these SNe in Appendix A. While half of our sample are SNe Ia-91T-like, strengthening the established connection between SNe Ia-CSM and SNe Ia-91T (Leloudas et al., 2015), we report the first detection of late-time MIR brightenings in a normal SN Ia, in a super-Chandra SN Ia,777The super-Chandra SNe Ia SN 2012dn (Nagao et al., 2017) and SN 2022pul (Siebert et al., 2024) have previously reported MIR excesses, but the lack of a detected MIR brightening are consistent with SN light echoes or new dust formation in cooling ejecta. and the second detection of an MIR excess in a SN Iax. We show a comparison of their MIR light curves in Figure 1, and the individual optical and MIR light curves of these events in Figure 2. Compared to normal SNe Ia, our newly identified events have significant MIR excesses reaching luminosities of Vega mag, lasting hundreds to thousands of days.
@@ 2.3 Multi-wavelength follow-up observations
We performed optical spectroscopic and radio follow-up of two of the youngest and closest events888We did not follow up SN 2021jun due to its proximity to its host nucleus, which would make faint object spectroscopy challenging.: SN 2020mvp () and SN 2022bbt (). Optical spectroscopic observations were carried out using a Fast Turnaround Program on the Gemini North telescope (Program ID GN-2024B-FT-102; PI: Mo), while radio observations were carried out using the Karl G. Jansky Very Large Array (VLA) under the Director’s Discretionary Program (VLA/24A-498; PI: Nayana AJ). In addition, we obtained NIR spectroscopy of SN 2020yex and NIR imaging of SN 2017fra using the Magellan Baade telescope, and NIR imaging of SN 2021jun and SN 2020mvp with WINTER (Lourie et al., 2020; Frostig et al., 2024). Further details about the observations and data reduction can be found in Appendix B. We do not detect photometric or spectroscopic signatures of CSM interaction in any of our optical/NIR observations while the VLA observations also did not detect a radio counterpart. We discuss the implications of these non-detections in Section 3.3.
@@ 3 Origin of the MIR emission
We have identified a sample of six SNe Ia showing clear evidence for late-time MIR brightenings. While fading excess late-time MIR emission can be caused by a light echo from the SN or dust formation in the cooling ejecta, rebrightening in the MIR light curve indicates the presence of an internal energy source (i.e., high-energy radiation in the form of UV or X-ray photons) that is heating pre-existing or newly-formed dust (Fox et al., 2010, 2011). Given that known SNe Ia exhibiting delayed CSM interaction (see Figure 1) also show late-time MIR brightenings, these observations strongly suggest that the brightenings are powered by the onset of CSM interaction. The lack of spectroscopic signatures at peak-light indicate that this sample represents an overlooked population of SNe Ia undergoing delayed CSM interaction. In this section, we use the NEOWISE multi-band photometry to constrain the mass, temperature, and location of the emitting dust to infer the CSM properties. In addition to the six SNe, we also fit the NEOWISE data from SN 2015cp, for comparison to a well-studied event.
@@ 3.1 Dust modeling
where is the observed spectral flux density, is the blackbody radius, is the distance to the source, and is the blackbody temperature. Given that the dust is expected to not emit as a pure blackbody source and may be optically thin, the derived is a lower limit to the true radius of the dust shell. Assuming the dust to be optically thin, the dust mass and temperature are described by
where is the dust mass, is the dust size, is the mass density of the dust, is a dimensionless quantity describing the dust emissivity at frequency , is the Planck function, and is the dust temperature (Hildebrand et al., 1977). We assume a dust grain size of 0.1 , a composition of carbonaceous grains, a dust mass density of g cm-3 (Draine & Li, 2007), and we use the emissivity function for graphite dust from Draine & Li (2001). The evolution of our fit parameters is shown in Figure 3.

### v4

@@ Abstract
Type Ia supernovae arise from the thermonuclear explosions of white dwarfs in multiple star systems. A rare sub-class of SNe Ia exhibit signatures of interaction with circumstellar material (CSM), allowing for direct constraints on companion material. While most known events show evidence for dense nearby CSM identified via peak-light spectroscopy (as SNe Ia-CSM), targeted late-time searches have revealed a handful of cases exhibiting delayed CSM interaction with detached shells. Here, we present the first all-sky search for late CSM interaction in SNe Ia using a new image-subtraction pipeline for mid-infrared data from the NEOWISE space telescope. Analyzing a sample of SNe Ia, we report evidence for late-time mid-infrared brightening in five previously overlooked events spanning sub-types SNe Iax, SNe Ia-91T and super-Chandra SNe Ia. Our systematic search doubles the known sample, and suggests that % of SNe Ia exhibit mid-infrared signatures of delayed CSM interaction. The mid-infrared light curves ubiquitously indicate the presence of multiple (or extended) detached CSM shells located at cm, containing of dust, with some sources showing evidence for new dust formation, possibly within the cold, dense shell of the ejecta. We do not detect interaction signatures in spectroscopic and radio follow-up; however, the limits are largely consistent with previously confirmed events given the sensitivity and observation phase. Our results highlight that CSM interaction is more prevalent than previously estimated from optical and ultraviolet searches, and that mid-infrared synoptic surveys provide a unique window into this phenomenon.
@@ 1 Introduction
Supporting the diverse nature of progenitors, a rare sub-class of SNe Ia—known as SNe Ia-CSM ( of SNe Ia; Silverman et al. 2013a; Sharma et al. 2023)—exhibit strong interaction of the expanding ejecta111We note that interaction with very nearby CSM can also be detected as early-time excesses in optical light curves or high velocity features (e.g., Noebauer et al. 2016; Mulligan & Wheeler 2017; Moriya et al. 2023); however, we do not discuss them here since existing observations do not conclusively rule out other possibilities (e.g. Childress et al. 2014; Deckers et al. 2022). with dense nearby circumstellar material (CSM). These events are typically identified via strong narrow emission lines (commonly H) in spectroscopy acquired at or near peak light (Sharma et al., 2023), similar to Type IIn supernovae (Smith, 2017). This H-rich material has been ubiquitously attributed to mass supplied by a non-degenerate companion, allowing for direct constraints on its composition and configuration. The proximity ( cm) and large mass ( M⊙) of the CSM have been explained via binary progenitors such as a WD merging with the core of an asymptotic giant branch (AGB) star inside (or shortly after the ejection) of the common envelope (i.e. the core-degenerate scenario; Tutukov et al. 1992; Livio & Riess 2003; Kashi & Soker 2011; Ablimit 2021) or even as a result of dynamical mass transfer from a main sequence star (Han & Podsiadlowski, 2006). Since systematic searches (Silverman et al., 2013a; Sharma et al., 2023) have relied on peak-light spectroscopy or peculiar undulations in early-time light curves to identify candidates for SNe Ia-CSM, these are biased against SNe Ia with distant CSM shells where signatures of interaction would only be detectable at later phases (Sharma et al., 2023; Phillips et al., 2024).
@@ 2.1 MIR behavior of known SNe Ia with late CSM interaction
Using data from the infrared Spitzer Space Telescope (Werner et al., 2004), Fox & Filippenko (2013) reported luminous, brightening late-time MIR emission in SNe Ia-CSM 2002ic and 2005gj, suggesting renewed shock interaction with CSM shells. Subsequently, Szalai et al. (2019), Szalai et al. (2021) and Sharma et al. (2023) also confirmed consistently bright late-time MIR emission in larger samples of SNe Ia-CSM based on Spitzer and NEOWISE observations, largely consistent with pre-existing dust heated by ongoing CSM interaction. A similar, luminous MIR excess was also reported in the case of the nearby SN Iax 2014dt (Fox et al., 2016; Jencson et al., 2019), although its origin has been debated (as arising from circumstellar dust or possibly a bound remnant; Foley et al. 2016). Using a comprehensive spectral sequence and NEOWISE observations, Wang et al. (2024) provided evidence for multiple dust shells as well as new late-time dust formation in the SN Ia-CSM 2018evt. To understand if MIR searches can be fruitful to look for delayed CSM interaction in SNe Ia, we first investigate previously known events in NEOWISE data. The NEOWISE survey’s all-sky coverage, MIR photometric bands (m, m), long baseline ( yr), regular cadence ( month), and sensitivity ( AB mag) are ideal attributes for such a search. When combined with a new difference photometry pipeline (De et al., in prep) with unWISE coadded images (Lang, 2014; Meisner et al., 2018) and a ZOGY-based subtraction algorithm (Zackay et al., 2016; De et al., 2020) adept at recovering faint transients on bright hosts, this MIR dataset has already proven to be powerful for core-collapse SNe (Myers et al., 2024).
NEOWISE observations of PTF11kx begin yrs after peak light, exhibiting luminous emission detected for two more years (also previously reported with Spitzer data in Graham et al. 2017). Similarly, the NEOWISE analysis of SN 2015cp reveals a long-lasting MIR light curve333Thévenot et al. (2021a) previously also reported the NEOWISE detection of SN 2015cp.. We show a comparison of these light curves with SNe Ia-CSM in Figure 1, and their individual light curves in Figure 5. Kool et al. (2023) reported luminous MIR emission from SN 2020eyj that exhibited delayed interaction with a He-rich CSM shell. In contrast, NEOWISE light curves of the sub-luminous SNe Ia which show late-time H emission (SNe 2016jae, 2018cqj, and 2018fhw), as well as the candidates identified from late-time optical photometry (Terwel et al., 2024) (SNe 2018grt, 2019ldf, and 2020tfc) do not show luminous MIR emission comparable to the other events. We discuss possible reasons in Section 3.3. As shown in Figure 1, many SNe Ia-CSM and SNe Ia with delayed CSM interaction show luminous late-time MIR emission, distinct from normal SNe Ia, making MIR searches a promising avenue for identifying this phenomenon.
@@ 2.2 Systematic search in NEOWISE for missed events
We produced NEOWISE MIR difference imaging light curves for Type Ia SNe from 2011 to 2022444The 2022 cutoff is governed by the availability of unWISE data products. reported to the IAU Transient Name Server555https://www.wis-tns.org. These light curves were visually inspected to search for late-time MIR emission associated with the SN. We filter the light curves by i) rejecting candidates with detections prior to the optical onset of the SN (to remove contamination from active galactic nuclei activity) or obvious subtraction artifacts; and ii) requiring at least two epochs of NEOWISE detections ( yr) and rising emission in at least one band, to conclusively rule out excess MIR emission arising purely from new dust formation or a SN light echo without added power from shock-interaction, and to eliminate detections deemed to only be associated with photospheric emission from the optical SN. Most of the identified candidates were known SNe Ia-CSM; we do not discuss them here as their light curves have been presented in Sharma et al. (2023). Table 2 summarizes the cuts at each filtering stage. Notably, these selection criteria resulted in the identification of five candidates which exhibit luminous late-time MIR emission despite not being previously associated with CSM interaction.666SN 2020mvp (Tonry et al., 2020) also emerged as a candidate, but further astrometric inspection of the NEOWISE imaging revealed that the MIR emission was likely from temporally coincident nuclear activity.
@@
The candidates span multiple Type Ia subclasses: SN 2017fra (originally classified as a SN Ia-91T), SN 2017hyn (Ia-91T), SN 2020yex (Ia-91T), SN 2021jun (Iax), and SN 2022bbt (super-Chandra Ia)777Thévenot et al. (2021a) and Thévenot et al. (2021b) have previously noted MIR detections for SN 2017hyn, SN 2017fra, and SN 2020yex; Jiang et al. (2021) previously analyzed NEOWISE observations of SN 2017fra in their search for MIR flares in nearby galaxies.. We summarize these SNe in Table 1, further describe each in Appendix A, and show reference, science, and difference image cutouts in Figure 6. While half of our sample are SNe Ia-91T-like, strengthening the established connection between SNe Ia-CSM and SNe Ia-91T (Leloudas et al., 2015), we report the first detection of late-time MIR brightenings in a super-Chandra SN Ia,888The super-Chandra SNe Ia SN 2012dn (Nagao et al., 2017) and SN 2022pul (Siebert et al., 2024) have previously reported MIR excesses, but the lack of a detected MIR brightening are consistent with SN light echoes or new dust formation in cooling ejecta. and the second detection of an MIR excess in a SN Iax. We show a comparison of their MIR light curves in Figure 1, and the individual optical and MIR light curves of these events in Figure 2. Compared to normal SNe Ia, our newly identified events have significant MIR excesses reaching luminosities of Vega mag, lasting hundreds to thousands of days. Of note is the MIR light curve of the SN Iax 2014dt, which shows similar behavior to but is much less luminous than the rest of the interacting SNe Ia; its late MIR rise has been suggested to arise from newly formed dust, but could also be attributable to pre-existing CSM or a bound remnant (Fox et al., 2016; Foley et al., 2016; Jencson et al., 2019).
Name | Type | Host morphology | Redshift | Distance | Peak W2 | Integrated W2 |
[Mpc] | AB magnitude | energy [erg] | ||||
SN 2017fra | Ia-91T | Interacting spiral | 0.03 | 136 | 16.8 | |
SN 2017hyn | Ia-91T | Edge-on disk | 0.053 | 244 | 18.5 | |
SN 2020yex | Ia-91T | Blue spheroidal | 0.087 | 425 | 19.3 | |
SN 2021jun | Iax | Blue spheroidal | 0.040 | 183 | 19.3 | |
SN 2022bbt | Ia-SC | Face-on spiral | 0.049 | 225 | 19.5 |
@@ 2.3 Multi-wavelength follow-up observations
We performed optical spectroscopic follow-up of two of the youngest events999We did not follow up SN 2021jun due to its proximity to its host nucleus, which would make faint object spectroscopy challenging.: SN 2020yex () and SN 2022bbt (). We also performed radio follow-up of SN 2022bbt. Optical spectroscopic observations were carried out using a Fast Turnaround Program on the Gemini North and Gemini South telescopes (Program IDs GN-2024B-FT-102/GS-2024B-FT-102; PI: Mo), while radio observations were carried out using the Karl G. Jansky Very Large Array (VLA) under the Director’s Discretionary Program (VLA/24A-498; PI: Nayana AJ). In addition, we obtained NIR spectroscopy of SN 2020yex and NIR imaging of SN 2017fra using the Magellan Baade telescope, and NIR imaging of SN 2021jun with WINTER (Lourie et al., 2020; Frostig et al., 2024). Further details about the observations and data reduction can be found in Appendix B. We do not detect photometric or spectroscopic signatures of CSM interaction in any of our optical/NIR observations while the VLA observations also did not detect a radio counterpart. We discuss the implications of these non-detections in Section 3.3.
@@ 3 Origin of the MIR emission
We have identified a sample of five SNe Ia showing clear evidence for late-time MIR brightenings. While fading excess late-time MIR emission can be caused by a light echo from the SN or dust formation in the cooling ejecta, rebrightening in the MIR light curve indicates the presence of an internal energy source (i.e., high-energy radiation in the form of UV or X-ray photons) that is heating pre-existing or newly-formed dust (Fox et al., 2010, 2011). The lack of observed UV or X-ray emission from these events is unsurprising given the non-existence of all sky UV/X-ray surveys and the expected luminosity of the shock interaction radiation. For the observed peak W2 MIR luminosity of erg s-1, the equivalent peak X-ray flux for sources at Mpc would be erg cm-2 s-1, which is far below the detection threshold of existing X-ray all-sky surveys ( erg cm-2 s-1 for MAXI/GSC, Hiroi et al. 2013; erg cm-2 s-1 for Swift/BAT, Krimm et al. 2013), and may be detectable only with pointed X-ray follow-up (which does not exist for these sources). Furthermore, targeted X-ray observations of other SNe Ia have had limited success due to the rarity (one known event in SN 2012ca; Bochenek et al. 2018) and low luminosity ( erg s-1 for normal SNe Ia from Margutti et al. 2014; erg s-1 for SNe Ia with delayed CSM interaction from Dwarkadas 2023, 2024) of X-ray emission associated with SNe Ia.
Given that known SNe Ia exhibiting delayed CSM interaction (see Figure 1) also show late-time MIR brightenings, these observations strongly suggest that the brightenings are powered by the onset of CSM interaction, where the shock of the high-velocity SN ejecta impacting existing CSM creates high-energy radiation that is absorbed and reemitted by dust. The integrated MIR energies shown in Table 1 span approximately 1-10% of the erg in kinetic energy released in a SN Ia due to the efficient conversion of kinetic energy to radiative energy during the shock; these shocks can convert up to of the total kinetic energy into radiative energy for interacting core-collapse SNe with similar CSM interaction (Smith, 2017). The lack of spectroscopic signatures or optical light curve abnormalities at peak-light indicate that this sample represents an overlooked population of SNe Ia undergoing delayed CSM interaction. In this section, we use the NEOWISE multi-band photometry to constrain the mass, temperature, and location of the emitting dust to infer the CSM properties. We assume a spherical configuration for the dust for simplicity, though our limited observations cannot rule out complex geometries which may affect the light curve or inferred parameters only up to order unity corrections. In addition to the five SNe, we also fit the NEOWISE data from SN 2015cp, for comparison to a well-studied event.
@@ 3.1 Dust modeling
where is the observed spectral flux density, is the blackbody radius, is the distance to the source, and is the blackbody temperature. Given that the dust is not expected to emit as a pure blackbody source and may be optically thin, the derived is a lower limit to the true radius of the dust shell. Assuming the dust to be optically thin, the dust mass and temperature are described by
where is the dust mass, is the dust size, is the mass density of the dust, is a dimensionless quantity describing the dust emissivity at frequency for dust size , is the Planck function, and is the dust temperature (Hildebrand et al., 1977). We assume a dust grain size of 0.1 m, a composition of carbonaceous grains, a dust mass density of g cm-3, and we use the emissivity function for graphite dust from Draine & Li (2001). Fits with different dust composition (silicate vs graphite), grain size (0.01 m vs 0.1 m), and dust mass density (1 g cm-3 vs 2.2 g cm-3 from Draine & Li 2007 vs 5 g cm-3) resulted in similar results, with dust masses within factors of a few and consistent dust temperatures. Since this fit assumes optically thin dust, and we lack observations at longer wavelengths which may help better constrain the dust, these are lower limits on dust masses. The evolution of our fit parameters is shown in Figure 3.